7. If the frequency of an allele d is 0.25 in a migrant population and 0.5 in The recipient population, and if the migration rate is 0.1, what is the frequency of d in the recipient population after one generation of migration?
Answer #1

Answer: 0.475 (As shown below)

Change in allele frequency in the recipient population is:

ΔPc = m( Po − Pr) = 0.1(0.25 - 0.5) = 0.1 (- 0.25) = - 0.025

where, ΔPc is the change in allele frequency in the recipent population,

Pr is the allele frequency in the recipient population

Po is the allele frequency in the original population

Therefore allele new frequency in the recipient population = Pc = Pr + ΔPc = 0.5 + (- 0.025) = 0.475